a.     The following information is from the national income accounts for country X

Y = C+I+G+(X-M)

C = 20+0.8Yd

T = 30

G = 22

X = 20

M = 4+0.3Y

Yd = Y-T

I=30

  1. informs the above model, list all the endogenous and exogenous variables
  2. Determine the equilibrium values for all the endogenous variables   
1
Expert's answer
2021-11-07T19:40:25-0500

(1) The endogenous are: taxes, exports, government expenditure, investments

Endogenous are: consumption and imports

(2) Y = C+I+G+(X-M)

"Y = 20 +0.8Yd + 30 +22 +(20 -(4 +0.3Y)) \\

Y = 20 +0.8(Y-30) + 30 +22 + (20 -4 -0.3Y) \\

Y = 20 + 0.8Y -24 +30 +22 +16 -0.3Y \\

Y-0.8Y +0.3Y =64 \\

0.5Y = 64 \\

Y = 128 \\

C = 20 +0.8(Y-30) \\

C = 20 +0.8(128-30) = 98.4 \\

M = 4+0.3(128) = 42.4"